# On the Application of Model Predictive Control to a Weighted Coverage Path Planning Problem

### v1

@@ Abstract
This paper considers the application of Model Predictive Control (MPC) to a weighted coverage path planning (WCPP) problem. The problem appears in a wide range of practical applications, such as search and rescue (SAR) missions. The basic setup is that one (or multiple) agents can move around a given search space and collect rewards from a given spatial distribution. Unlike an artificial potential field, each reward can only be collected once. In contrast to a Traveling Salesman Problem (TSP), the agent moves in a continuous space. Moreover, he is not obliged to cover all locations and/or may return to previously visited locations. The WCPP problem is tackled by a new Model Predictive Control (MPC) formulation with so-called Coverage Constraints (CCs). It is shown that the solution becomes more effective if the solver is initialized with a TSP-based heuristic. With and without this initialization, the proposed MPC approach clearly outperforms a naive MPC formulation, as demonstrated in a small simulation study.
@@ I Introduction
Many path planning applications in robotics try to find an optimal path while maximizing some sort of reward. The reward is commonly related to the proximity to given reference value, guiding the system via an artificial potential field (AFP), or keeping it away from unsafe areas via barrier functions. Model Predictive Control (MPC) casts this problem into a numerical optimization program, with an objective function and constraints. Over the past years, it has become a standard approach for path planning, due to its natural handling of general reward functions, system dynamics, and input and state constraints.
Coverage Path Planning (CPP) aims for the system to cover an entire area of the state space, or as large a part of it as possible. Problem instances appear in many robotic applications, such autonomous lawn mowers, vacuum cleaners, agricultural robots, and arial / underwater drones used for inspection or surveillance. From the perspective of common MPC-based path planning, this means a uniform objective function. However, for CPP, the objective function changes dynamically in the sense that the reward at an already visited position, and some radius around it, drops to zero.
Efficient motion patterns or policies are commonly used for CPP, including a boustrophedon (snake-like) path or straight driving with random reflection angles when hitting boundaries [1]. The main principle is, clearly, to minimize any overlaps in the track of the robot.
This paper considers the extended problem of Weighted Coverage Path Planning (WCPP). The main difference to the CPP is that the objective function is not uniform, but weighted by a coverage priority. Problem instances of this also appear frequently, e.g., in search and rescue (SAR) or surveillance missions, where the goal is to find or detect a target whose probability of presence on a map is not uniformly distributed. Correspondingly, the motivational example for this paper is an unmanned aerial vehicle (UAV) with the task of finding a missing person in a given area as fast as possible. The coverage priority is indicated by a probability map, which serves as the reward function and represents a belief distribution integrating all available information at the current time, e.g., from sensor measurements, the observation of eye witnesses, or human behavioral models in the given map [2]. For practical applications, a probabilistic model could be used to dynamically update the belief distribution based on all available information, and thus guide the planning of the mission [3].
I-A Existing Literature
CPP is about finding a path that covers all specified points or an entire area or region of interest while avoiding obstacles [4]. A comprehensive overview of CPP algorithms for robotic applications is given in [5]. The article covers both classical and heuristic-based algorithms. These two categories include a whole variety of basic approaches, such as AFPs, greedy search and graph search algorithms, and bio-inspired approaches, such as genetic algorithms or ant colony optimization.
@@ I-A Existing Literature
A key point in the discussion in [5] is that the resulting paths should be as smooth as possible. Namely, the avoidance of sharp turns in the path prevent premature wear of the robot’s components and it increases the efficiency, especially in UAV applications, where it advantageous for the average speed of the UAV to fly straight ahead or in smooth curves. In this spirit, the approach in [6] proposes a path search algorithm using ant colony optimization based on the Lin-Kernighan heuristic, followed by a smoothing step using a customized approach based on Fourier series. The resulting dynamically smooth trajectory is then fed to the UAV, which is operated by an MPC-based controller.
@@
Numerical optimization has become increasingly popular for path planning over the recent years, due to the availability of more powerful hardware and increasingly efficient solvers. MPC, in particular, has been successfully employed for collision-free path planning for autonomous road vehicles [7] and for UAVs [8]. MPC has been used in combination with AFPs [9, 10], and also for CPP with obstacles using a mixed-integer linear programming (MILP) formulation [11]. The area is covered by a uniform grid, where each cell defines a single way point. The way points are subsequently represented as discrete decision variables within the optimization problem, where the objective is to cover the maximum number of (equally weighted) way points.
@@
Related problems to the CPP include the Traveling Salesman Problem (TSP) and its variants, most notably the Orienteering Problem (OP) [12]. In contrast to the TSP, where all vertices must be visited and the goal is to minimize the traveled distance, the OP is concerned with maximizing the total reward collected within a limited time frame, without necessarily visiting all the vertices [13]. Both of these problems are known to be NP-hard [14]. Yet it is desirable, for many applications, to extend them further by including the dynamics of an agent in the problem formulation [15]. Recently, a mathematical framework to tackle this problem using techniques from optimal control has been proposed [15, 16]. In addition to the dynamics of an agent, this formulation also accounts for the movement of the sensors, i.e., in this case, a camera. However, the approach relies on nonsmooth calculus and considers only discrete regions of interest.
@@ II-A Finding a trajectory with MPC
We consider a discrete-time optimal control problem of finding the optimal sequence of control inputs $u_{0},u_{1},…,u_{N-1}\in\mathbb{R}^{m}$, driving the states $x_{0},x_{1},…,x_{N}\in\mathbb{R}^{n}$ of a dynamical system $x_{k+1}=f(x_{k},u_{k})$ for $k=0,1,…,N-1$, where $N$ is the horizon length.
@@
The constraints (II-A) and (II-A) keep the states and inputs in the set of admissible states (the region of the map) $\mathbb{X}$ and in the set of admissible inputs (accelerations) $\mathbb{U}$, respectively. The constraint (II-A) sets the initial state of the agent.
@@
Note that the dynamics (II-A) in the MPC formulation are generally arbitrary. For the implementation, simple linear dynamics $f(x_{k},u_{k})=Ax_{k}+Bu_{k}$ are assumed, in the form of a double integrator:
@@ II-B Finding an initial solution with a TSP based heuristic
The presented algorithm is based on Gaussian mixture models (GMMs). They can approximate any probability density function [17] and have also been used before to prioritize search subregions in a probability map [18]. Note that it is not necessary to consider the reward in (1) as a probability density function. However, we make the additional assumptions that the reward is always positive and that the total collectable reward is finite:
@@ A Gaussian mixture with $n$ components is a function given by
To feed the travelling salesman tour from the as an initial solution to the MPC problem (II-A), the tour is discretized using the maximum velocity as the step size. Figure 4 shows the result of running the MPC as the closed-loop controller, given the tour from Figure 3 as the initial solution.
@@ III Results
The methods described in Section II have been implemented in Julia, using the JuMP toolkit [22] for defining optimization problems. As the underlying solver, HiGHS [23] has been used for solving mixed-integer problems (TSP), and Ipopt [24] for non-linear continuous optimization problems (MPC). The parameters of the objective function (II-A) have been selected as $c_{1}=1,c_{2}=1000,c_{3}=100$. All experiments have been performed on a machine with an Intel i5-1350P processor.

### v2

@@ Abstract
This paper considers the application of Model Predictive Control (MPC) to a weighted coverage path planning (WCPP) problem. The problem appears in a wide range of practical applications, including search and rescue (SAR) missions. The basic setup is that one (or multiple) agents can move around a given search space and collect rewards from a given spatial distribution. Unlike an artificial potential field, each reward can only be collected once. In contrast to a Traveling Salesman Problem (TSP), the agent moves in a continuous space. Moreover, he is not obliged to cover all locations and/or may return to previously visited locations. The WCPP problem is tackled by a new Model Predictive Control (MPC) formulation with so-called Coverage Constraints (CCs). It is shown that the solution becomes more effective if the solver is initialized with a TSP-based heuristic. With and without this initialization, the proposed MPC approach clearly outperforms a naive MPC formulation, as demonstrated in a small simulation study.
@@ I Introduction
State-of-the-art methodologies in robotic path planning typically seek to determine an optimal trajectory, maximizing a reward function given the system’s dynamic capabilities. The reward function is often a weighted sum of terms for particular purposes, e.g., proximity-based guidance by Artificial Potential Fields (APF) or enforcing safety boundaries via barrier functions. Model Predictive Control (MPC) fully encompasses this framework by formulating the problem as a numerical optimization program. Over the past decade, MPC has become a standard approach for many path planning applications, owing to its inherent capacity to effectively integrate general objective functions, complex system dynamics, and constraints on the system’s inputs and states.
Coverage Path Planning (CPP) is characterized by the mandate for the system to traverse, or cover, the entire area of the designated state space, instead of moving to a target area or point. Instances of this problem are prevalent in diverse robotic domains, including autonomous agricultural systems, automated cleaning devices, and aerial or underwater drones employed for inspection or surveillance. Motion planning approaches are typically based on pre-defined patterns or simple policies, including, but not limited to, boustrophedon (snake-like) patterns or random-walk strategies involving boundary reflections [1]. The core principle underpinning these approaches is the minimization of spatial and temporal overlaps in the system’s trajectory.
Weighted Coverage Path Planning (WCPP) is an extension of CPP. The primary difference is the reward is not uniformly distributed, as for the CPP problem. Instead, it is weighted by an explicit coverage priority. Example applications include
  • (i)
    a mobile irrigation system, which may aim to prioritize areas based on their moisture level, preferring areas with particularly dry soil;
  • (ii)
    a cleaning robot, whose goal may be to collect as much dirt as possible, hence using a coverage priority based on the actual distribution of dirt;
  • (iii)
    an aerial or underwater inspection or surveillance system, whose priority distribution is related to the probability of detecting a fault or finding a target in a given area or on a given map.
CPP problems usually assume a static environment, such as the cleaning of a room. WCPP problems, on the other hand, often appear in dynamic settings, such as the surveillance for moving targets. Clearly, accounting for dynamically changing environment adds a extra level of complexity, and it is also possible to consider WCPP for a static, yet not uniform coverage priority.
This paper focusses on the static WCPP. It is motivated by search and rescue (SAR) missions, e.g., in the Wadden Sea, using an Unmanned Aerial Vehicle (UAV). The goal is to locate a lost person as quickly as possible. Thus the coverage priority is given by a belief distribution of the person’s current location, as a function over the search area. This so-called probability map may be based on information about the last known location of the person, a behavioral model, the environmental conditions, and the elapsed time [3]. More details on the generation of the probability map and a comparison of search methods can be found in a recent publication [2].
This paper proposes an MPC formulation for the static WCPP. From the perspective of MPC, the priority function, or probability map, can be considered as a reward function, or a negative cost function. The major difference to conventional MPC is that each reward can only be collected once, because upon entering a certain position, the related priority drops to zero. In the case of the SAR mission, for example, the UAV is supposed to plan a trajectory to collect as much of the probability mass as possible, instead of hovering above the point with the highest probability value.
I-A Literature Review
Many algorithms have been proposed for CPP in different environments with and without obstacles [4, 5]. Algorithms can be divided into randomized and deterministic, and they are based on techniques such as AFPs, greedy search, graph search algorithms, and bio-inspired approaches, like genetic algorithms or ant colony optimization.
@@ I-A Literature Review
A key point for some applications is that the resulting paths should be smooth [5]. In fact, some robots are not able to turn on the spot. Sharp turns may also slow the robot down or lead to premature wear of the robot’s components. For example, in [6], a search algorithm based on ant colony optimization with the Lin-Kernighan heuristic, is combined with a path smoothing using Fourier series. The resulting trajectory is then fed to the UAV, which is operated by a controller based on MPC.
@@
MPC itself has been successfully employed for collision-free path planning for autonomous road vehicles [7] as well as UAVs [8]. It has been combined with with AFPs [9, 10], and it has also been employed previously for CPP with obstacles using a mixed-integer linear programming (MILP) formulation [11]. To this end, the area is covered by a uniform grid, where each cell defines a single way point. The way points are subsequently represented as discrete decision variables within the optimization problem, where the objective is to cover the maximum number of (equally weighted) way points.
@@
Related to CPP and WCPP is the Traveling Salesman Problem (TSP) and its variants, most notably the Orienteering Problem (OP) [12]. In contrast to the TSP, where the salesman has to visit all vertices of the graph and the goal is to minimize the traveled distance, the OP is concerned with maximizing the total reward collected within a limited time frame, without necessarily visiting all the vertices [13]. Both problems are known to be NP-hard [14]. Yet it is desirable, for many applications, to extend them by considering a spacial map instead of a graph, and including the dynamics of an agent in the problem formulation [15]. A mathematical framework to tackle this problem using techniques from optimal control has been recently proposed [15, 16]. In addition to the dynamics of an agent, this formulation also accounts for the movement of a sensor, in this case, a camera. However, the approach relies on non-smooth calculus and considers only discrete regions of interest.
@@ II-A Finding a trajectory with MPC
We consider a discrete-time optimal control problem of finding the optimal sequence of control inputs $u_{0},u_{1},…,u_{N-1}\in\mathbb{R}^{m}$, driving the states $x_{0},x_{1},…,x_{N}\in\mathbb{R}^{n}$ of a dynamical system $x_{k+1}=f(x_{k},u_{k})$ for k=0,1,,N1k=0,1,…,N-1, where $N$ is the horizon length.
Denote by $\mathbf{u}_{0:N-1}:=\{u_{0},u_{1},…,u_{N-1}\}$ the sequence of control inputs and by $\mathbf{x}_{0:N}:=\{x_{0},x_{1},…,x_{N}\}$ the corresponding sequence of states. Let $r(x_{k})$ be the reward at state $x_{k}$. Then the (non-linear) optimal control problem is given by
$\underset{\displaystyle\mathbf{x}_{0:N},\mathbf{u}_{0:N-1}}{\mathrm{min}}\quad J(\mathbf{x}_{0:N},\mathbf{u}_{0:N-1})\,$, (2a)
@@
$x_{k+1}=f(x_{k},u_{k})\quad$ $\forall\;k=0,…,N-1,$ (2b)
$x_{k}\in\mathbb{X}$ $\forall\;k=0,…,N,$ (2c)
$u_{k}\in\mathbb{U}$ $\forall\;k=0,…,N-1,$ (2d)
$x_{0}=x(0)\;.$ (2e)
The objective function $J(\mathbf{x}_{0:N},\mathbf{u}_{0:N-1})$ contains the control input costs and the state rewards:
@@
The constraints (2c) and (2d) keep the states and inputs in the set of admissible states (the region of the map) $\mathbb{X}$ and in the set of admissible inputs (accelerations) $\mathbb{U}$, respectively. The constraint (2e) sets the initial state of the agent.
@@
Note that the dynamics (2b) in the MPC formulation are generally arbitrary. For the implementation, simple linear dynamics $f(x_{k},u_{k})=Ax_{k}+Bu_{k}$ are assumed, in the form of a double integrator:
@@ II-B Finding an initial solution with a TSP based heuristic
The presented algorithm is based on Gaussian mixture models (GMMs). They can approximate any probability density function [17] and have also been used before to prioritize search subregions, e.g., based on a probability map [18, 2]. Note that it is not necessary to consider the reward in (1) as a probability density function. However, we make the additional assumptions that the reward is always positive and that the total collectable reward is finite:
@@ A Gaussian mixture with $n$ components is a function given by
To feed the travelling salesman tour from the as an initial solution to the MPC problem (2), the tour is discretized using the maximum velocity as the step size. Figure 4 shows the result of running the MPC as the closed-loop controller, given the tour from Figure 3 as the initial solution.
@@ III Results
The methods described in Section II have been implemented in Julia, using the JuMP toolkit [22] for defining optimization problems. As the underlying solver, HiGHS [23] has been used for solving mixed-integer problems (TSP), and Ipopt [24] for non-linear continuous optimization problems (MPC). The parameters of the objective function (2a) have been selected as $c_{1}=1,c_{2}=1000,c_{3}=100$. All experiments have been performed on a machine with an Intel i5-1350P processor.